What is ΔGo (in kJ) at 647 K for the following reaction? PbO(g) + CO2(g) → PbCO3(s) PbO: ΔHfo = -219.0 kJ/mol and So = 66.5 J/K mol) PbCO3(s): ΔHfo = -699.1 kJ/mol and So = 131.0 J/K mol) CO2: ΔHfo = -393.5 kJ/mol and So = 213.6 J/K mol)

At what temperature (in K) does the above reaction become spontaneous?
Answer #1

Hrxn = Hprod - Hreact = -699.1 - (-219.0 + -393.5 ) = -86.6 kJ/mol

Srxn = Sprod - Sreact = 131.0 - (213.6 +66.5 ) = -149.1 J/molK

then

G = H - TS

G < 0 so it is spontanous

H - TS < 0

H/S < T

T > H/S

T > -86600/(-149.1 ) = 580.818 K

T must be higher than 580.818 K